OpenStudy (anonymous):

The perimeter of a rectangle is 42 centimeters. The length of the rectangle can be represented by x+4 and it's width can be represented by 2x-7. What are the dimensions of the rectangle in centimeters?

OpenStudy (anonymous):

The length is 12 and the width is 9 12 + 12 = 24 9 + 9 = 18 18 + 24 = 42